For each of the reactions, calculate the mass (in grams) of the product formed when 15.42 g of the underlined reactant completely reacts. Assume that there is more than enough of the other reactant.
2Sr(s)–––––+O2(g)→2SrO(s)
Express your answer using four significant figures.
1
Expert's answer
2020-04-06T11:51:40-0400

m(SrO)= m(Sr)*2*Mr(SrO)/2*Mr(Sr) = 15.42*2*104/2*88=18.22 g

Answer: 18.22 g